find three consecutive integers such that four times the sum of all three is two times the product of the larger two
Answer by nerdybill(7384)

Let x = 1st consecutive integer
x+1 = 2nd
x+2 = 3rd
.
4(x + x+1 + x+2) = 2(x+1)(x+2)
4(3x + 3) = 2(x^2+3x+2)
12x + 12 = 2x^2+6x+4
0 = 2x^2-6x-8
0 = x^2-3x-4
0 = (x-4)(x+1)
.
x = {4,-1}
.
Solution: 4,5,6